OpenStudy (trojanpoem):

A wide section cylinder container of radius 10 cm with a side gab at its bottom ( gab cross section area = 1 cm^2). Water was poured into it at a fixed rate1.4 * 10^-4 m^3/s Find the time needed to empty the container if water pouring stopped.

ganeshie8 (ganeshie8):

Cross section area the cylinder is \(\pi*10^2~cm^2\). Notice that this area is much greater than the gab area which is \(1~cm^2\)

ganeshie8 (ganeshie8):

Hey is that the entire question ?

OpenStudy (trojanpoem):

@ganeshie8 , I was away. Nah, The rest : the speed of water at the gab: \[v = \sqrt{2gh}\] g is the gravity, h is the height from the water level to the gab Pouring ratio = Q = AV A cross section , V speed, Q pouring ratio Figure : 

OpenStudy (trojanpoem):

the pouring rate is constant.

OpenStudy (loser66):

Assume the tank is empty at the beginning. We have 2 periods: 1) Water in - water out = water stays. Hence after t1 seconds, the tank will be full and we stop pouring the water in. At that time, period 2 starts. 2) just find out the time to empty the tank. Hence, if my assuming is correct, then we just assume the tank is full at the beginning to start calculating period 2. I don't understand the question since we don't know where to start, the empty tank? the full tank? when we stop pouring the water in? It is over my head, :) Can't wait to see the solution. :)

OpenStudy (loser66):

@dan815 @freckles @ikram002p @ganeshie8 please.

OpenStudy (trojanpoem):

You start after the tank was totally full.

OpenStudy (trojanpoem):

Ops, I forgot, the water height when water pouring stopped = 0.1 m xD was on the figure.

OpenStudy (ikram002p):

I can't jelp now. But if u have time maybe @irishboy123

ganeshie8 (ganeshie8):

\[v = \sqrt{2gh} = \dfrac{dx}{dt}\]